A sealed container with negligible coefficient of volumetric expansion contains helium When it is heated from 300K to 600K the average KE of helium atom is

Grade:12

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallvi

Average K.E of a molecule is given as

K.E =3/2 kT

chenge in K.E  = 3/2 kT - 3/2kT1

=3/2k (T - T1)

=3/2 * 1.38 * 10-23 ( 600 -300)

=6.21 * 10-21  J

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.
 All the best.
 
Regards,
Askiitians Experts
Badiuddin

Abhinav M
6 years ago
The answer its doubled as KE is directly proportional to T . ie. =3/2xKTAs the KE here is directly proprtinal to Temperature a change in temperature will effect change in KE linearly Hope u got ur answer!